### v1

@@ Abstract.
While anonymity networks such as Tor provide invaluable privacy guarantees to society, they also enable all kinds of criminal activities. Consequently, many blameless citizens shy away from protecting their privacy using such technology for the fear of being associated with criminals. To grasp the potential for alternative privacy protection for those users, we design Seldom, an anonymity network with integrated selective deanonymization that disincentivizes criminal activity. Seldom enables law enforcement agencies to selectively access otherwise anonymized identities of misbehaving users, while providing technical guarantees preventing these access rights from being misused. Seldom further ensures translucency, as each access request is approved by a trustworthy consortium of impartial entities and eventually disclosed to the public (without interfering with ongoing investigations). To demonstrate Seldom’s feasibility and applicability, we base our implementation on Tor, the most widely used anonymity network. Our evaluation indicates minimal latency, processing, and bandwidth overheads compared to Tor, while Seldom’s main costs stem from storing flow records and encrypted identities. With at most of storage required in total to retain the encrypted identifiers of a Tor-sized network for two years, Seldom provides a practical and deployable technical solution to the inherent problem of criminal activities in anonymity networks. As such, Seldom sheds new light on the potentials and limitations when integrating selective deanonymization into anonymity networks.
@@ 1. Introduction
Tor (Dingledine et al., 2004) and other anonymity networks provide invaluable services to, e.g., journalists, whistleblowers, and militaries, who are among its millions of daily users (The Tor Project, Inc., 2024). One may argue that Tor anonymizes too well, as criminals misuse Tor and established a hotspot of highly illegal activities, such as sharing Child Sexual Abuse Material (CSAM) (Nurmi et al., 2024). Reports showing that over of Ahmia.fi searches seek CSAM (Nurmi et al., 2024) and over of marketplace listings are related to illegal drugs (Hiramoto and Tsuchiya, 2023) are indeed worrying. Collectively, researchers and some governmental agencies may accept this fact as a necessarily evil for the irreplaceable service Tor provides to society.
@@
On an individual level, however, people do have morale concerns regarding such misuse of Tor. Studies highlight that Tor users are concerned about illegal activities within the network (Winter et al., 2018; Gallagher et al., 2017). Moreover, a study suggests that of US citizens “would like to do more” to improve their privacy, while only of participants have used anonymity networks such as Tor in the past (Rainie and Madden, 2015). Consequently, there exists a large untapped user base for anonymity services. We argue that parts of these users, potentially with only a passing understanding of Tor, shy away from it due to its close association with illegal activities they want to stay away from. After all, each user indirectly supports all Tor activities—by increasing everyone’s anonymity (Dingledine and Mathewson, 2006)—and (exit) relay operators become active targets of criminal investigations.
@@
With this paper, we strive to investigate the potential advantages and limitations of an alternative anonymity network that offers integrated selective deanonymization of suspicious users (Seldom). Seldom empowers Law Enforcement Agencies to effectively prosecute crime by investigating suspicious activities within the anonymity network with due cause and under public oversight, while technically limiting these capabilities to prevent abuse of power. Seldom thus provides those who do not want to be associated with Tor an alternative to adequately protect their anonymity on the Internet. As the provided anonymity grows with the number of users (Dingledine and Mathewson, 2006), wide acceptance of Seldom can offer strong anonymity even when LEA can deanonymize selected suspicious identities based on publicly agreed-upon rules.
@@
Seldom realizes its functionality to deanonymize a small number of communication flows based on a novel oblivious authentication protocol that links all outbound traffic to threshold-encrypted user identities. A trustworthy consortium of impartial parties then decides about LEA’ deanonymization requests for individual communication flows. To provide public transparency without impeding ongoing investigations, we propose an immutable but delayed public release of all activities via a translucent ledger. Comparing the performance of Seldom to Tor, we measure only an imperceivable latency and processing overhead. Seldom would thus offer a Tor-like experience without the downside of supporting crime when merely using the network. If significantly outgrowing Tor, Seldom could even reduce network latency, and thus address one of the major drawbacks of onion routing.
@@
Contributions. To provide a thorough understanding of the potential of selective deanonymization in anonymity networks, we make the following contributions in this paper:
@@
We devise an oblivious authentication protocol that links each outgoing communication flow to a unique threshold-encrypted identity of the anonymized client and we integrate this protocol into our anonymity network Seldom.
@@
We design a translucent deanonymization procedure that only reveals the client identity to the requesting LEA.
@@
Our evaluation shows that Seldom has an imperceivable performance impact on the users’ experience and imposes only well-manageable data storage requirements on centralized databases.
@@
Finally, we discuss the challenges, limitations, and implications of deploying an anonymity network such as Seldom with exceptional deanonymization capabilities.
@@ 2. Acceptable Exceptional Access
Providing data access to LEA in just circumstances is often referred to as exceptional access. Any system providing such exceptional access is justly viewed critically. It is thus vital to technologically protect individual privacy by requiring regulation, transparency, and accountability, as LEA could otherwise undermine that privacy. However, to effectively uncover crime, LEA need good oversight and actionability. To combine these conflicting requirements, we lay out the prerequisites to successfully strike a balance in the safety-privacy dilemma. Without prioritization due to the contrary perspectives on this matter, we identify the following essential aspects of acceptable exceptional access by LEA.
@@
Translucency. LEA should operate transparently and inform the public of their surveillance activities. However, timely access to such data also helps criminals evade surveillance. Hence, translucency, which allows for critical information to remain temporarily classified, is needed.
Accountability. If a LEA abuses its given privileges, its exceptional access privileges must be revocable as a reaction by the general public.
@@
Limited User Overhead. When using a system with exceptional access, citizens should not be unnecessarily affected by overhead stemming from enabling this access.
Overall, these prerequisites ought to ensure that surveillance of digital communication is (i) limited to exceptional cases, (ii) effective in banishing crime from the network, and (iii) acceptable for the citizens who shall receive protection.
@@ 3. Related Work
In the following, we give an overview of previous directions to enable exceptional access. We notice that related work has evolved and considerably broadened its scope since the initial proposals during the 1990s crypto wars.
Partial Key Escrow. Early works focused on (partial) key escrow, which involves the government storing (parts of) all used encryption keys (Denning and Branstad, 1996; Bellare and Goldwasser, 1997). Hence, the government would be able to retrieve keys and brute-force partial keys if needed. However, such systems require strong trust in key holders; even the protection of partial keys decays over time as computation power rises (Abelson et al., 2015). Similarly, the idea of obliviously escrowing encryption keys at random peers (Blaze, 1996) was quickly abandoned due to overhead and trust concerns. Recent work looks at the use of distributed ledgers as trusted parties for key escrow (Green et al., 2021; Fetzer et al., 2023). These systems, however, require significant user cooperation and cannot currently be implemented efficiently. Overall, it is also unclear how people with criminal intent would be motivated to use such key escrow systems.
Translucent Cryptography. Orthogonally to key escrow schemes, translucent cryptography (Bellare and Rivest, 1999) is based on a variant of oblivious transfers (Rabin, 1981) called Non-Interactive Fractional Oblivious Transfers (NFOTs). In this scheme, any message also holds an Non-Interactive Fractional Oblivious Transfer (NFOT), which allows an LEA to decrypt wiretapped messages with a predefined probability. However, at least one communication partner must remain honest in this scheme. Further, evaluating a wiretapped NFOT leaves no evidence, which prevents transparency or accountability. Finally, a purely probabilistic per-message scheme does not hold up with today’s communication patterns and surveillance capabilities. Namely, being able to decrypt only a tiny fraction of messages still allows one to peek into a large fraction of conversations but prohibits gaining all relevant information from individual rightfully investigated individuals.
@@
Lawful Device Access. Another proposal considers the design of special storage hardware that can be decrypted based on prolonged physical access, e.g., hours to days (Savage, 2018). LEA would then have to prove continual physical access before being able to read the storage. This way, transparency, accountability, and rate-limited access are established by requiring traditional seizures of hardware. Namely, this approach could have provided a legitimate framework for accessing the device at the center of the controversy around the San Bernadino shooting (Thielman, 2016). However, the roll-out of such technology would be slow due to its reliance on the wide adoption of specialized storage hardware, and it does not consider encrypted containers stored on such hardware.
@@
Privacy-preserving Content Scanning. Orthogonally, client-side scanning wants to identify illegal material such as CSAM on clients’ devices. In this vein, privacy-preserving content scanning attempts to realize client-side scanning while respecting the user’s privacy (Kulshrestha and Mayer, 2021; Bartusek et al., 2023). However, current approaches suffer from limitations, e.g., the uncertainty of whether detection remains restricted to harmful material and will not be abused for censorship or surveillance of political opponents (Abelson et al., 2024). Indeed, recent analysis shows how the underlying perceptual hashing algorithms to compare images are susceptible to adversarial images, i.e., images in the database that look like CSAM to humans but which hashes collide with other secretly tracked content (Jain et al., 2022, 2023; Hooda et al., 2024).
@@
Deanonymization in Onion Routing. A wide range of approaches strives to deanonymize activities in anonymity networks by performing traffic analysis such as website fingerprinting on the connection between client and entry relay, e.g., to derive which website a certain client likely visited. To this end, different approaches claim to identify visited websites based on traditional machine learning (Panchenko et al., 2011; Wang et al., 2014; Hayes and Danezis, 2016; Dyer et al., 2012; Cai et al., 2012; Panchenko et al., 2016), deep learning (Rimmer et al., 2018; Sirinam et al., 2019; Shen et al., 2023; Rahman et al., 2020; Bhat et al., 2019), or generative artificial intelligence (Oh et al., 2021). However, besides various issues with website fingerprinting itself such as unrealistic or oversimplified assumptions (Juarez et al., 2014) as well as limited scalability (Panchenko et al., 2016; Cherubin et al., 2022), these approaches neither provide strict correctness guarantees nor allow to deanonymize clients when observing activity related to the server-side, e.g., who accessed a certain website. To the best of our knowledge, BackRef (Backes et al., 2014) is the only proposal to provide such functionality by enabling relays to iteratively prove that they are not the origin of suspicious traffic. BackRef (Backes et al., 2014) is promising but creates significant storage overhead for all relays and requires them to prove their innocence, which becomes problematic if relay operators cannot provide evidence, e.g., due to data loss. Finally, BackRef (Backes et al., 2014) is neither transparent in terms of scale nor purpose of requests.
@@
In summary, past work on exceptional access aims to give tools to LEA to protect public safety. However, this comes with unavoidable privacy concessions that are usually regarded critically as such services can provide a slippery slope towards mass surveillance. In contrast, we are the first to propose to use exceptional access capabilities to increase privacy by providing a new service to an untapped user base currently browsing the Internet without anonymity.
@@ 4. High-Level Overview of Seldom
We now give a high-level overview of the design of Seldom before diving into its technical details in the following Sections 5 and 6. First, we begin however by laying out the threat model considered in this work.
4.1. Threat Model
We consider the same threat model for Seldom that is applied to Tor (Dingledine et al., 2004). Hence, we assume that there exists no global passive adversary who could observe and match all incoming and outgoing traffic of each relay. Instead, the considered adversaries can observe a fraction of all Internet traffic and also operate a portion of relays. These direct passive observation are also one way for LEA to identify critical traffic on the Internet that may require deanonymization. Alternatively, LEA can request (usually through a warrant) traffic metadata, e.g., IP addresses, timestamps, from Internet Service Providers or website operators.
@@ 4.2. Design Overview
Seldom is an anonymity network based on onion routing that is extended with selective deanonymization capabilities according to our prerequisites formulated in Section 2. We design Seldom by extending Tor’s design. Figure 1 provides a high-level overview of Seldom.
Clients use Seldom just like Tor, but Seldom securely stores data to retrieve the clients’ identities in a database under shared control of a consortium of impartial parties. This data allows an LEA observing suspicious traffic leaving Seldom to file a deanonymization request with the consortium, who can jointly reveal the client’s identity to the LEA.
@@
Immutably logging all requests and decisions is crucial for maintaining transparency towards the general public. However, Seldom has to ensure that criminals under investigation do not gain an advantage via this channel. To address this issue, we introduce a translucent ledger (cf. Section 6.2) that is maintained by the consortium and allows for a delayed full disclosure while providing a general overview of ongoing activities in real time. With this form of public oversight, citizens can notice abuses of power within Seldom and abandon the network in cases of government overreach. In case of a successful deanonymization, users also only lose exactly the anonymity gained through the use of the network, i.e., their browsing is as safe or unsafe as if no anonymity network was ever used.

### v2

@@ Abstract.
While anonymity networks such as Tor provide invaluable privacy guarantees to society, they also enable all kinds of criminal activities. Consequently, many blameless citizens shy away from protecting their privacy using such technology for fear of being associated with criminals. To grasp the potential for alternative privacy protection for those users, we design Seldom, an anonymity network with integrated selective deanonymization that disincentivizes criminal activity. Seldom enables law enforcement agencies to selectively access otherwise anonymized identities of misbehaving users while providing technical guarantees preventing these access rights from being misused. Seldom further ensures translucency, as each access request is approved by a trustworthy consortium of impartial entities and eventually disclosed to the public (without interfering with ongoing investigations). To demonstrate Seldom’s feasibility and applicability, we base our implementation on Tor, the most widely used anonymity network. Our evaluation indicates minimal latency, processing, and bandwidth overheads compared to Tor; Seldom’s main costs stem from storing flow records and encrypted identities. With at most of storage required in total to retain the encrypted identifiers of a Tor-sized network for two years, Seldom provides a practical and deployable technical solution to the inherent problem of criminal activities in anonymity networks. As such, Seldom sheds new light on the potentials and limitations when integrating selective deanonymization into anonymity networks.
@@ 1. Introduction
Tor (Dingledine et al., 2004) and other anonymity networks provide invaluable services to, e.g., journalists, whistleblowers, and militaries, who are among its millions of daily users (The Tor Project, Inc., 2024). One may argue that Tor anonymizes too well, as criminals misuse Tor to establish a hotspot of highly illegal activities, such as sharing Child Sexual Abuse Material (CSAM) (Nurmi et al., 2024). Reports showing that over of Ahmia.fi search sessions seek CSAM (Nurmi et al., 2024) and over of marketplace listings are related to illegal drugs (Hiramoto and Tsuchiya, 2023) underpin the need for reevaluating the societal downsides of unconditionally anonymous Internet use.
@@
Collectively, privacy advocates, such as researchers and some governmental agencies, may accept this fact as a necessary evil for the irreplaceable service Tor provides to society. On an individual level, however, people do have moral concerns regarding such misuse of Tor. Studies highlight that Tor users are concerned about illegal activities within the network (Winter et al., 2018; Gallagher et al., 2017). Moreover, a study suggests that of US citizens “would like to do more” to improve their privacy, while only of participants have used anonymity networks such as Tor in the past (Rainie and Madden, 2015). Consequently, there exists a potentially large untapped user base for anonymity services. We argue that some of these users, potentially with only a passing understanding of Tor, shy away from it due to its close association with illegal activities they intend to stay away from. After all, each user indirectly supports all Tor activities—by increasing everyone’s anonymity (Dingledine and Mathewson, 2006)—and (exit) relay operators become active targets of criminal investigations.
@@
With this paper, we strive to investigate the potential advantages and limitations of an alternative anonymity network that offers integrated selective deanonymization of suspicious users (Seldom). Seldom empowers Law Enforcement Agencies to effectively prosecute crime by investigating suspicious activities within the anonymity network only with acknowledged due cause and under public oversight, while technically limiting these capabilities to prevent abuse of power at the same time. Seldom thus provides those who do not want to be associated with Tor an alternative to adequately protect their anonymity on the Internet. As the provided anonymity grows with the number of users (Dingledine and Mathewson, 2006), wide acceptance of Seldom can offer strong anonymity even when LEA can deanonymize selected suspicious identities based on publicly agreed-upon rules.
@@
Seldomrealizes its functionality to deanonymize a fraction of communication flows based on a novel oblivious authentication protocol that links all outbound traffic to threshold-encrypted user identities. A trustworthy consortium of impartial parties then decides about LEA’ deanonymization requests for individual communication flows. To provide public transparency without impeding ongoing investigations, we propose an immutable but delayed public release of all activities via a translucent ledger. Comparing the performance of Seldom to Tor, we measure only an imperceptible latency and processing overhead. Seldom would thus offer a Tor-like experience without the downside of supporting crime when merely using the network. If significantly outgrowing Tor, Seldom could even reduce network latency and thus address one of the major drawbacks of onion routing.
@@
Contributions. To provide a thorough understanding of the potential and implications of selective deanonymization in anonymity networks, we make the following contributions:
@@
We devise an oblivious authentication protocol that links each outgoing communication flow to a unique threshold-encrypted identity of the anonymized client, and we integrate this protocol into our anonymity network Seldom.
@@
We design a translucent deanonymization process that only reveals the client identity to the requesting LEA.
@@
Our evaluation shows that Seldom has an imperceptible performance impact on the users’ experience and imposes only well-manageable data storage requirements on centralized databases.
@@
Finally, we discuss the challenges, limitations, and implications of deploying an anonymity network with exceptional deanonymization capabilities.
@@ 2. Acceptable Exceptional Access
Providing data access to an LEA in legitimate circumstances is often referred to as exceptional access. Any system providing exceptional access naturally faces public scrutiny due to its potential for abuse. It is thus vital to technologically protect individual privacy against LEA by enforcing regulation, transparency, and accountability. However, LEA need to be able to identify unfolding crimes and act accordingly. To combine these conflicting requirements, we lay out the prerequisites to successfully strike a balance in the safety-privacy dilemma. Overall, we identify the following aspects to be essential for a widely acceptable form of exceptional access:
@@
Translucency. LEA should operate transparently and inform the public of targeted surveillance activities. However, timely access to such data also helps criminals evade surveillance. Hence, translucency, which allows for critical information to remain temporarily classified, is needed.
Accountability. If an LEA abuses its given privileges, its exceptional access privileges must be revocable in a way that is transparent to the system’s users.
@@
Limited User Overhead. When using a system offering exceptional access, citizens should not be unnecessarily affected by overhead stemming from enabling this access.
Overall, these essential prerequisites ought to ensure that surveillance of digital communication (i) is limited to exceptional cases, (ii) is effective in prosecuting crime on the network and thereby disincentivizing it, and (iii) generates wide societal acceptance by strengthening public safety.
@@ 3. Related Work
In the following, we give an overview of previous approaches to enabling exceptional access. We notice that related work has evolved and considerably broadened its scope since the initial proposals during the 1990s crypto wars.
Partial Key Escrow. Early works focused on (partial) key escrow, which involves the government maintaining a global store of (parts of) all used encryption keys (Denning and Branstad, 1996; Bellare and Goldwasser, 1997). Hence, the government would be able to retrieve keys and brute-force partial keys if needed. However, such systems require strong trust in key holders; even the protection of partial keys decays over time as computation power rises (Abelson et al., 2015). Similarly, the idea of obliviously escrowing encryption keys at random peers (Blaze, 1996) was quickly abandoned due to overhead and trust concerns. Recent work looks at the use of distributed ledgers as trusted parties for key escrow (Green et al., 2021; Fetzer et al., 2023). These systems, however, require active adoption from users, while it remains unclear how users with criminal intent would be motivated to use such key escrow systems.
Translucent Cryptography. Orthogonally to key escrow schemes, translucent cryptography (Bellare and Rivest, 1999) is based on a variant of oblivious transfers (Rabin, 1981) called Non-Interactive Fractional Oblivious Transfers (NFOTs). In this scheme, any message also holds an Non-Interactive Fractional Oblivious Transfer (NFOT), which allows an LEA to decrypt wiretapped messages with a predefined probability. However, at least one communication partner must remain honest in this scheme. Further, evaluating a wiretapped NFOT leaves no evidence, which prevents transparency or accountability. Finally, a purely probabilistic per-message scheme does not hold up with today’s communication patterns and surveillance capabilities. Namely, being able to decrypt only a tiny fraction of messages still allows one to peek into a large fraction of conversations but prohibits gaining all relevant information from any rightfully investigated individual.
@@
Lawful Device Access. Another proposal considers the design of special storage hardware that can be decrypted based on prolonged physical access, e.g., hours to days (Savage, 2018). LEA would then have to prove continual physical access before being able to read the storage. This way, transparency, accountability, and rate-limited access are established by requiring traditional seizures of hardware. Namely, this approach could have provided a legitimate framework for accessing the device at the center of the controversy around the San Bernardino shooting (Thielman, 2016). However, the roll-out of such technology would be slow due to its reliance on the wide adoption of specialized storage hardware, and it does not consider encrypted containers stored on such hardware.
@@
Privacy-preserving Content Scanning. Orthogonally, client-side scanning wants to identify illegal material such as CSAM on clients’ devices. In this vein, privacy-preserving content scanning attempts to realize client-side scanning while respecting the user’s privacy (Kulshrestha and Mayer, 2021; Bartusek et al., 2023). However, current approaches suffer from limitations, e.g., the uncertainty of whether detection remains restricted to harmful material and will not be abused for censorship or surveillance of political opponents (Abelson et al., 2024). Indeed, recent analysis shows how the underlying perceptual hashing algorithms to compare images are susceptible to adversarial images, i.e., images in the database that look like CSAM to humans but have hashes that collide with other secretly tracked content (Jain et al., 2022, 2023; Hooda et al., 2024).
@@
Deanonymization in Onion Routing. A wide range of approaches strives to enable deanonymizations of activities in anonymity networks through traffic analysis, such as website fingerprinting on the connection between client and entry relay, e.g., to derive which website a client likely visited. To this end, different approaches claim to identify visited websites based on traditional machine learning (Panchenko et al., 2011; Wang et al., 2014; Hayes and Danezis, 2016; Dyer et al., 2012; Cai et al., 2012; Panchenko et al., 2016), deep learning (Rimmer et al., 2018; Sirinam et al., 2019; Shen et al., 2023; Rahman et al., 2020; Bhat et al., 2019), or generative artificial intelligence (Oh et al., 2021). However, besides various issues with website fingerprinting itself, such as unrealistic or oversimplified assumptions (Juarez et al., 2014) or limited scalability (Panchenko et al., 2016; Cherubin et al., 2022), these approaches neither provide correctness guarantees nor allow deanonymizing clients when observing activity related to the server side, e.g., who accessed a certain website. To the best of our knowledge, BackRef (Backes et al., 2014) is the only proposal to provide such functionality by enabling relays to iteratively prove that they are not the origin of suspicious traffic. BackRef (Backes et al., 2014), however, creates significant storage overhead for all relays and requires them to prove their innocence, which becomes problematic if relay operators cannot provide evidence, e.g., due to data loss. Finally, BackRef (Backes et al., 2014) is neither transparent in terms of scale nor purpose of requests.
@@
Past work on exceptional access thus aims to provide LEA with tools to intrude on an individual’s privacy to ultimately protect public safety. In contrast, we are the first to propose using exceptional access capabilities to increase privacy by providing a new service to an untapped user base currently browsing the Internet without anonymity.
@@ 4. High-Level Overview of Seldom
In this section, we give a high-level overview of Seldom’s design and the assumed threat model before we describe Seldom’s technical details in greater depth in the following Sections 5 and 6.
@@ 4.1. Design Overview
Seldomis an anonymity network based on onion routing that is extended with exceptional access capabilities according to our prerequisites formulated in Section 2. We design Seldom by extending Tor’s design. Figure 1 provides a high-level overview of Seldom’s architecture. Like Tor, the Seldom network uses an ensemble of directory authorities to manage the relays in the network and also supports bridge relays that do not need to be listed publicly.
Clients use Seldom just like Tor, but Seldom securely stores data related to the clients’ identities in a database under shared control of a consortium of impartial parties. This process allows an LEA observing suspicious Internet traffic leaving the Seldom network to file a substantiated deanonymization request with the consortium. It is the consortium that then jointly decides whether it should reveal the client’s identity to the requesting LEA, with the knowledge that this decision and the attached justification will be released to the public in the future.
@@
[…]
Immutably logging all requests and decisions is crucial for maintaining transparency toward the users. However, Seldom has to ensure that criminals under investigation do not gain an advantage via this channel. To address this issue, we introduce a translucent ledger (cf. Section 6.2) that is maintained by the consortium and allows for a delayed full disclosure while providing a general overview of ongoing activities in real time. With this form of public oversight, citizens can notice patterns of potential power abuses within Seldom and abandon the network in cases of government overreach. In case of a successful deanonymization, users also only lose exactly the anonymity gained through the use of the network, i.e., their browsing is as safe or unsafe as if no anonymity network was ever used.
4.2. Threat Model and Trust Assumptions
The threat model for Seldom and the trust assumptions toward the different entities involved are guided by the core goal of Seldom:
In the following, we first introduce our threat model before discussing the trust assumptions toward the different entities in the network.
4.2.1. Threat Model
We build on Tor’s threat model (Dingledine et al., 2004) and extend it to capture the deanonymization capabilities introduced by Seldom. Hence, we assume that there exists no global passive adversary who is capable of matching all incoming and outgoing traffic at every relay. Instead, we consider adversaries that can observe a fraction of all Internet traffic and possibly operate a portion of relays. This direct passive observation is one of the ways for LEA to identify critical traffic on the Internet that may require deanonymization. Alternatively, LEA can request (usually through a warrant) traffic metadata, e.g., IP addresses and timestamps, from Internet Service Providers or website operators. While a single person or entity can take on multiple roles in the network, it cannot take over a significant share of e.g., relay operators. An LEA can, for example, take on the role as client and relay operator in the network at the same time. Further, a government, as one entity, could even act as an LEA and a consortium member. With these capabilities, the goals of an adversary can be twofold. They can either evade or prevent deanonymizations to hide suspicious activities, frame an innocent client, or simply disrupt the network to hinder adoption. On the other hand, an adversary can also attack Seldom to conduct unauthorized deanonymizations.
4.2.2. Trust Assumptions about the Different Roles
In the following, we list the entities involved in the operation of Seldom and which trust assumptions are made about them.
Client. Anyone can operate one or multiple Seldom clients, but a client only has access to information about itself.
IP Certificate Authority. IP CAs are independent entities that have to be designated when Seldom is set up. Deanonymization results of Seldom can only be trusted by the requesting LEA to the degree they trust the CA.
Seldom Relay. Any entity can operate one or multiple relays. As for Tor, clients need to trust that no single entity controls a large fraction of all relays, as proper anonymization is not guaranteed otherwise. Relays are expected to follow the Seldom protocol and will be excluded from the network if misbehavior is detected.
Seldom Exit Relay. Exit relays are relays that are exposed to the Internet and thus the initial suspicion points for LEA. In Tor, this risk deters many relay operators. On the other hand, Seldom, with its automated deanonymization process, can better protect exit relays. In turn, exit relays are expected to be available for deanonymizations.
Directory Authority. Seldom, like Tor, has a set of highly trusted directory authorities that manage the list of relays.
Consortium. The members of the consortium must be carefully chosen, and only a small number of changes over time are expected. It must collectively be trusted to act honestly. For Seldom to function properly, the consortium thus has to be trusted by clients, relays, and LEA alike, while each individual may distrust different members within the consortium.
Database Operator. We assume that all entities trust the database operator(s). The involved risk is that they can see the metadata of all outgoing connections through flow records. As detailed later in Section 5.3 (Figure 3), database operators do not see which flows are associated with the same circuit. In practice, the database could be operated by the directory authorities, by governments to redistribute costs if the privacy implications are accepted, or distributed among multiple entities.